# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 5.5.2 Using Data More Efficiently
Self-supervised training enhances data efficiency in reward model training by using unannotated data to capture information about the task. This technique extends beyond the scope of pretraining methods, as discussed in the prior section, to incorporating concurrent auxiliary tasks to maximize the utility of available data. A prevalent technique, as applied by [2], [39], and [180], involves adding self-supervised losses to enhance representation learning for rewards. [2] implement a contrastive task where the reward network differentiates between observations that are consistent between multiple modalities and those that are not, blending this with preference learning loss and behavioral cloning. [39] add multiple auxiliary tasks such as inverse and forward dynamics modeling, temporal distance prediction, and variational autoencoder training. Similarly, [180] use the self-predictive representations technique [229] to learn state representations that encode environmental dynamics, enabling a linear model to anticipate successor states, thereby forming an efficient basis for preference learning and significantly boosting sample efficiency. However, auxiliary losses for better representation learning are not the only approach to leverage self-supervised training. An alternate approach by [260] involves identifying important states using attention weights from a world model transformer and state importance estimates based on a preference predicting transformer. These estimates can aid credit assignment for observed preferences, further optimizing the training process.
@@
Semi-supervised training, blending labeled and unlabeled data, can leverage the unlabeled data to glean information about the task and the environment. This is most commonly done by generating pseudo-labels for the unlabeled data, either by leveraging model predictions or by making assumptions. The first approach is utilized by [44] and [285], which use generative models and GAN-based methods to mimic human preferences. Similarly, [203] expand their data set with high-confidence unlabeled samples based on the preference predictor’s evaluations. The second strategy, making assumptions to augment data, is showcased by [293]. They generate preference data by assuming that (i) human-written examples are better than model-written examples, (ii) human-written and model-written examples are indistinguishable amongst themselves, and (iii) generations of later model iterations are better than those of earlier ones.
@@
Data augmentation focuses on creating additional examples from existing labeled data. Temporal augmentation is particularly effective in RLHF, involving trajectory data. This is exemplified by [40, 203] who base their augmentation on the premise that preferences for complete trajectories can be extrapolated to cropped segments, allowing the generation of multiple derivative pairs from a single labeled trajectory pair. [203] additionally explore state modifications, such as random re-scaling and Gaussian noise addition, finding temporal cropping to be the most effective, with noise sometimes negatively impacting performance. In a similar vein, [258] focus on augmenting trajectories by concentrating on changing elements in observations and perturbing the other parts, based on the premise that movement indicates importance in image-based observations. Complementing these methods, [2] employ augmentation by randomly altering instructions and language responses, thus creating artificial examples of non-preferred behavior. These diverse data augmentation methods collectively enhance the training data set, contributing to the increased robustness and efficacy of RLHF models.
@@ 5.5.3 Gathering Better Data
While we are not aware of many works in this area, one possible approach involves steering the agent’s exploration towards regions of the state space where human feedback would be most beneficial. [162] implement this by employing intrinsic motivation, driven by the estimated uncertainty of the reward model, to guide the agent’s exploration. This highlights the potential of not just using data more efficiently but also generating data in a more targeted manner.
@@ 6 Theory
The field of RLHF has recently made some progress in terms of theoretical results, which we will discuss in this section. First, we consider the contributions where the goal is to learn a provable (near) optimal policy both in an online and offline fashion or even in a way that falls in between. Then, we discuss and highlight recent contributions related to different theoretical aspects of RLHF, such as its relation to the standard reward-based RL. Tables 5 and 6 provide a concise overview of the results for the online or offline policy learning setting. Here, $\mathcal{N}_{\mathcal{F}}(\epsilon,d)$ denotes the $\epsilon$-covering number of a set $\mathcal{F}$ under some metric $d$ [^3] The $\epsilon$-covering number is the minimum integer $N$ such that there exists a subset $\mathcal{F}^{\prime}\subset\mathcal{F}$ with $|\mathcal{F}^{\prime}|=N$, and for any $f\in\mathcal{F}$, there exists some $f^{\prime}\in\mathcal{F}^{\prime}$ satisfying $d(f,f^{\prime})\leq\epsilon$.. It is worth mentioning that all works have two standard assumptions, namely that the reward function is bounded and that the ground-truth reward, human feedback model, or transition dynamic is an element of the considered model space, respectively.
@@ 6.1 Policy Learning
In the literature focusing on theoretical results, there is a distinction (similar to the distinction made in standard RL) between an offline and online setting. In the former, learning is based on a given fixed data set, usually previously collected through an interaction with the environment. In contrast, in the online environment, one interacts directly with the environment to learn from real-time feedback and continuously updates one’s strategies based on the feedback received, allowing the agent to learn and adapt as it engages with the environment.
@@
[196]
Leveraging Posterior Sampling from dueling bandits
Linear link function, tabular MDP
Bayes regret minimization w.r.t. optimal policy based on trajectory comparison feedback
Asymptotic regret rate:
@@
[225]
Leveraging MaxInP from contextual dueling bandits
Logistic link function, tabular MDP, linear rewards & $d$-dimensional feature embedding of trajectories
Expected regret minimization w.r.t. optimal policy based on trajectory comparison feedback
Transition dynamics:
@@ 2. Unknown
[52]
Leveraging MaxInP; general function approximation
General human feedback model class $\mathcal{F}_{\mathbb{T}}$ and general transition dynamic class $\mathcal{F}_{\mathbb{P}}$ with finite $l_{2}$-norm $\rho$-Eluder dimension $d_{\mathbb{T}}^{(2)}(\rho)$ and $d_{\mathbb{P}}^{(2)}(\rho)$, respectively
High probability regret minimization w.r.t. optimal policy based on trajectory comparison feedback
$\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{P}}(\tfrac{1}{T})HT\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{P}}}\left(\tfrac{1}{T},d\right)\right)}\right)\newline +\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{T}}(\tfrac{1}{T})T\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{T}}}\left(\tfrac{1}{T},d\right)\right)}\right)$
@@
[280]
Dynamic programming, policy search, $(\epsilon,\delta)$-PAC black-box dueling bandit algorithm and simulator
Uniform dependence of policy preference probabilities on value function differences, tabular MDP, $(\epsilon,\delta)$-PAC dueling bandit algorithm with $\Psi(K,\varepsilon,\delta)\varepsilon^{-\alpha}$ sample complexity for $K$ arms
$(\epsilon,\delta)$-PAC for optimal policy based on trajectory comparison feedback
Simulator step bound
@@
[280]
Similar to PPS, instead of simulator using an auxiliary synthetic reward function
Same as PPS and stochastic triangle inequality of trajectory comparisons preferences
$(\epsilon,\delta)$-PAC for optimal policy based on trajectory comparison feedback
Step complexity bound
@@
[145]
Optimistic least-squares value iteration, maximum information gain, value iteration based on pessimistic expected value function estimation
Binary rewards for state-action pairs based on human response model $f\in\mathcal{F}_{H}$ with bounded noise $\Delta>0$, compliant and tabular/linear MDP with dimension $d$
$(\epsilon,\delta)$-PAC for optimal policy based on binary state-action reward feedback
Tabular MDP:
@@
[274]
Least-squares value iteration with perturbed state-action-wise reward model
General differentiable link function $\Phi$, linear MDP, linear rewards with $d$-dimensional feature embedding of trajectories
Expected regret minimization w.r.t. optimal policy and/or low trajectory comparison feedback complexity steered by $\epsilon\in[0,1]$
Expected regret bound:
@@
The first work dealing with the question of theoretical guarantees for learning an optimal policy from trajectory comparison feedback (see Section 3.2) in an online manner is by [196]. It laid the foundation for a paradigm subsequently embraced by many subsequent research endeavors: Adapting learning algorithms from the dueling or preference-based bandit literature [284, 247, 21] to the underlying situation with additional states. The preference-based bandit problem can be viewed as a preference-based RL problem with one state, so state transition dynamics must be considered accordingly for a fruitful adaptation. It is worth mentioning that [122] used a quite similar idea before for feedback in the form of corrections (see Section 3.2) by resorting to the coactive learning setting [233]. Assuming the existence of a ground-truth context-trajectory scoring function and that the user’s feedback is informative, the Preference Perceptron algorithm by [233] is used and analyzed in terms of its cumulative regret.
@@
[196] suggest the Dueling Posterior Sampling (DPS), which is an adaptation of the self-sparring algorithm [246]. It takes a Bayesian perspective on the problem and defines a Dirichlet prior on the dynamics and a Gaussian prior on the rewards that are subsequently updated, while the trajectories to be compared by the human labeler are chosen based on their (posterior) probability of being optimal [^4] The latter probability is assessed by posterior sampling; a commonly used technique in the bandit literature used by so-called Thompson Sampling strategies, see [152] for more details.. Assuming a linear link function (see Section 5.1.4) as well as a tabular MDP, it is shown that DPS is (i) consistent, i.e., converges in distribution to the optimal policy, and (ii) achieves an asymptotic expected regret bound (see Table 5).
@@
[280] combine dynamic programming and policy search with a black-box preference-based bandit algorithm for each state to design routines that return an almost optimal (a.k.a. $\varepsilon$-optimal) policy with high probability [^5] This is a so-called PAC learning setting [254] in which the goal of finding the/an optimal object is relaxed to finding a “good enough” object, usually specified by some distance measure on the object domain. The first routine, called Preference-based Policy Search (PPS), requires access to a simulator, while the second routine, called Preference-based Exploration and Policy Search (PEPS), gets rid of this requirement by exploring the state space by means of an auxiliary synthetic reward function. By assuming that the probability of one policy dominating another policy is bounded uniformly over all states from below by a multiplicative of their value function, they show generic upper bounds for both routines on the number of pairwise trajectory comparisons (see Table 5). If these dominance probabilities have even more structural properties, such as fulfilling stochastic transitivity or stochastic triangle inequality (see [106, 21]), then these upper bounds can be further refined.
@@
A follow-up work by [225] assumes a feature embedding of trajectories that gives rise to a feature embedding of policies and adapts the MaxInP algorithm [224] for contextual dueling bandits by essentially viewing the policy embeddings as the contexts. More precisely, assuming a logistic link function (see Section 5.1.4), confidence sets for the expected scores of the policies are constructed based on the maximum likelihood estimate (MLE), and the two policies with the highest uncertainty in terms of maximal variance are used to sample a trajectory, respectively, to be compared. In this way, the logistic preference-based reinforcement learning (LPbRL) is derived and also extended to the case of unknown dynamics by taking the uncertainty regarding the dynamics into account when constructing the confidence sets. For both cases, i.e., known or unknown dynamics, upper bounds on the regret of LPbRL are shown (see Table 5).
@@
In contrast to previous work that all considers tabular MDPs, [52] consider the case of a general unknown human feedback model and unknown dynamics each from function classes with a finite Eluder dimension [^6] Roughly speaking, the Eluder dimension of a function class refers to the number of worst-case errors one must make to identify an unknown function from that class. [220]. They propose and analyze the Preference-based Optimistic Planning (PbOP) algorithm, which essentially follows a similar design as LPbRL but uses least-square estimates for the human feedback model and transitions dynamics along with confidence sets based on them. Moreover, they derive lower bounds for the regret of any learning algorithm by reducing the once-per-episode-feedback RL problem [49] to the PbRL problem. Finally, they extend their analysis to the case of $K$-wise comparisons, where one obtains all $\binom{K}{2}$ pairwise comparisons for $K$ many queried trajectories. In essence, the regret term coming from the human feedback model class improves by a factor of $\sqrt{K}$.

### v2

@@ 5.5.2 Using Data More Efficiently
Self-supervised training enhances data efficiency in reward model training by using unannotated data to capture information about the task. This technique extends beyond the scope of pretraining methods, as discussed in the prior section, to incorporating concurrent auxiliary tasks to maximize the utility of available data. A prevalent technique, as applied by Abramson et al. (2022), Brown et al. (2020), and Metcalf et al. (2023), involves adding self-supervised losses to enhance representation learning for rewards. Abramson et al. (2022) implement a contrastive task where the reward network differentiates between observations that are consistent between multiple modalities and those that are not, blending this with preference learning loss and behavioral cloning. Brown et al. (2020) add multiple auxiliary tasks such as inverse and forward dynamics modeling, temporal distance prediction, and variational autoencoder training. Similarly, Metcalf et al. (2023) use the self-predictive representations technique (Schwarzer et al., 2021) to learn state representations that encode environmental dynamics, enabling a linear model to anticipate successor states, thereby forming an efficient basis for preference learning and significantly boosting sample efficiency. However, auxiliary losses for better representation learning are not the only approach to leverage self-supervised training. An alternate approach by Verma & Metcalf (2024) involves identifying important states using attention weights from a world model transformer and state importance estimates based on a preference predicting transformer. These estimates can aid credit assignment for observed preferences, further optimizing the training process.
@@
Semi-supervised training, blending labeled and unlabeled data, can leverage the unlabeled data to glean information about the task and the environment. This is most commonly done by generating pseudo-labels for the unlabeled data, either by leveraging model predictions or by making assumptions. The first approach is utilized by Cao et al. (2021b) and Zhan et al. (2021), which use generative models and GAN-based methods to mimic human preferences. Similarly, Park et al. (2022) expand their data set with high-confidence unlabeled samples based on the preference predictor’s evaluations. The second strategy, making assumptions to augment data, is showcased by Zhou & Xu (2020). They generate preference data by assuming that (i) human-written examples are better than model-written examples, (ii) human-written and model-written examples are indistinguishable amongst themselves, and (iii) generations of later model iterations are better than those of earlier ones.
@@
Data augmentation focuses on creating additional examples from existing labeled data. Temporal augmentation is particularly effective in RLHF, involving trajectory data. This is exemplified by Brown et al. (2019) and Park et al. (2022) who base their augmentation on the premise that preferences for complete trajectories can be extrapolated to cropped segments, allowing the generation of multiple derivative pairs from a single labeled trajectory pair. Park et al. (2022) additionally explore state modifications, such as random re-scaling and Gaussian noise addition, finding temporal cropping to be the most effective, with noise sometimes negatively impacting performance. In a similar vein, Verma & Kambhampati (2023b) focus on augmenting trajectories by concentrating on changing elements in observations and perturbing the other parts, based on the premise that movement indicates importance in image-based observations. Complementing these methods, Abramson et al. (2022) employ augmentation by randomly altering instructions and language responses, thus creating artificial examples of non-preferred behavior. These diverse data augmentation methods collectively enhance the training data set, contributing to the increased robustness and efficacy of RLHF models.
Relatedly, Meta-Reward-Net (Liu et al., 2022) optimizes not only for the preference prediction accuracy of the learned reward function but also of the learned $Q$ function in an actor-critic RL algorithm. This is beneficial since it avoids the phenomenon of confirmation bias, where one learned model (in this case the $Q$ function) overfits to targets predicted by another model (the reward model). It is not strictly a data augmentation technique, but closely related in practice.
@@ 5.5.3 Gathering Better Data
While we are not aware of many works in this area, one possible approach involves steering the agent’s exploration towards regions of the state space where human feedback would be most beneficial. Liang et al. (2022) implement this by employing intrinsic motivation, driven by the estimated uncertainty of the reward model, to guide the agent’s exploration. This highlights the potential of not just using data more efficiently but also generating data in a more targeted manner.
6 Policy Learning
[…]
Figure 5: Highlighting the components of the RLHF loop discussed in this section.
After learning a reward model, or, more commonly, interleaved with reward model learning, the next step is to train a policy that maximizes the expected accumulated reward. This section will algorithms for policy learning, which can be categorized into two main techniques: adaptation of conventional RL algorithms and direct policy optimization (DPO).
6.1 Adaptation of RL Algorithms
Using the learned reward model, any standard RL algorithm (e.g., DQN, A3C, PPO, SAC) could potentially be applied to train a policy. However, in the setting of RLHF, this direct application may suffer from two issues: The non-stationary nature of the learned reward function in RLHF and the inaccuracy of intermediate reward models. We will discuss these issues and possible adaptations of RL algorithms to address them in the following.
Non-Stationary Rewards
RL algorithms are designed to learn a policy that maximizes the expected accumulated reward in an MDP framework, which assumes a stationary reward function. The RLHF setting violates that assumption by periodically updating the reward model, leading to a non-stationary reward function.
Various works have empirically demonstrated that conventional RL algorithms can be applied nonetheless, with little to no modification. Christiano et al. (2017) argue that policy-gradient methods are better suited for non-stationary reward functions compared to value-based methods. They and various follow-up works successfully apply policy-gradient methods without any modifications in this setting. This approach has been picked up for language-model fine-tuning as well (Ouyang et al., 2022).
Later works have shown that value-based methods (possibly in an actor-critic scheme) can also be effective in RLHF (Ibarz et al., 2018; Lee et al., 2021b; Park et al., 2022; Liu et al., 2022; Xue et al., 2023b). One trick to make value-based methods work is to use the reward model to relabel the experiences in the replay buffer whenever it is updated (Ibarz et al., 2018; Lee et al., 2021b). Similar to conventional RL, the use of such a replay buffer can greatly decrease the amount of environment interactions necessary for successful learning. As demonstrated by Gulcehre et al. (2023), the sample efficiency can be increased even further by using offline-RL techniques in a growing-batch RL setting, an offline-RL technique that iteratively increases the size of the dataset by policy rollouts while still being more sample-efficient than online RL.
In addition to the basic RL approaches, there are also some policy learning approaches tailored specifically for RLHF. Wu et al. (2023) propose a policy gradient algorithm, called Pairwise Proximal Policy Optimization (P3O), as an alternative to PPO, which avoids estimating the value function and at the same time is provably invariant with respect to equivalent rewards (unlike PPO). In a similar vein, Zhu et al. (2023b) replace the KL-regularization of PPO by means of a squared error term of the logarithmic probabilities resulting in a seemingly more stable RL learner.
Overoptimization of Approximate Rewards
Since the learned reward model, which is only an approximation of the true reward function, is used to train a policy, overoptimization (Gao et al., 2023) or reward hacking (Skalse et al., 2022) can happen. Section 5.2.5 discusses the interdependence of the reward model and the policy in more detail as well as possible improvements from the reward model side, while here we focus on how to adapt policy training to cope with possibly inaccurate rewards in general.
One approach is to regularize the policy so as not to diverge too much from human-given demonstrations. This is particularly common for language-model fine-tuning (Ouyang et al., 2022; Abramson et al., 2022), but Abramson et al. (2022) explores this for control as well. They found that this was important for some cases, in particular for deciding when to output language, but not for all. Going beyond KL-regularization, Moskovitz et al. (2024) investigate several techniques of constrained RL to only maximize rewards up to a threshold while avoiding excessive deviation from a pre-trained policy.
6.2 Framing RLHF for Generative Models as a Bandit Problem
So far, we have assumed that we ultimately want to solve a reinforcement learning problem represented by an MDP. However, especially with regard to the application of RLHF for the area of LLMs, there is now another simplified way of looking at the problem. Namely, as an instantiation of a (contextual) preference-based bandits problem (Bengs et al., 2021), which can of course be modeled by the more general case of a Markov decision process (MDP). In both cases, the focus is on the concept of tokens or rather sequences of tokens. However, in the MDP point of view, the state space ${\mathcal{S}}$ consists of all previous tokens and the prompt (represented as a sequence of tokens), while the action space ${\mathcal{A}}$ consists of all potential next tokens. A terminal state is often indicated here by the special token <eos> and trajectories are filled with this token until the maximum length $H$ is reached. Moreover, the transition function ${P}$ is degenerated (or deterministic) with being one only for the state that is the concatenation of the current state and the taken action. A (latent) reward is only received at the end of the trajectory giving rise to a sparse feedback scenario.
The (contextual) preference-based bandits view, on the other hand, naturally considers no state space and no transition function, but an action space consisting of all possible responses to a prompt (both represented as a sequence of tokens). Here, the prompt specifies the context for which at least two actions are executed and for which a qualitative comparison is observed as feedback. In bandit literature, this is also referred to as a “(multi-)duel”, coining the term dueling bandits. Thus, this point of view takes a trajectory-wise perspective, while the MDP point of view takes a token-wise perspective.
Note that the bandit formulation considers an entire episode (response in the LLM context) as an action with a single associated reward, resulting in sparse feedback. This is in contrast to the standard RLHF formulation as it is often used in control settings, where it is assumed that the reward of a trajectory is composed of the sum of the rewards of individual steps, which allows the optimizer to distribute rewards densely as best fits the data. On an intuitive level, this leads to state-action pairs that often occur in preferred trajectories to be highly rewarded, without necessarily putting all reward on the terminal actions. In practice, this can lead to nicely-shaped reward functions (Christiano et al., 2017), which cannot directly be achieved in the bandit setting. However, Chan et al. (2024) show how to take advantage of the predominantly used transformer architecture for the reward model in order to obtain a denser reward, even when assuming the bandit setting: More specifically, since the transformer architecture maintains attention weights in the last layer for each token, these can be used to attribute the overall reward signal to individual tokens.
The main appeal of the bandit formulation is that it does not require exploration of the environment’s dynamics, since they are deterministic. It therefore enables simpler policy learning approaches, such as DPO (Rafailov et al., 2023) or $\Psi$PO (Azar et al., 2023), discussed in the following section.
6.3 Direct Policy Optimization
The two-phase approach involving utility learning and policy optimization is not the only viable path to learning a policy from human feedback. While we have previously discussed the case in which we learn a reward function from observed preferences by assuming a human feedback model, an emerging branch of the literature is concerned with circumventing the reward-learning step and using preferences directly to address the actual RL problem. Concrete approaches are DPO (Rafailov et al., 2023), SLiC-HF (Zhao et al., 2023), OPPO (Kang et al., 2023), DPPO (An et al., 2023), PRO (Song et al., 2024), RSO (Liu et al., 2024b), or by formulating policy search as a zeroth-order optimization (Tang et al., 2024). Azar et al. (2023) introduce an objective called $\Psi$-preference optimization ($\Psi$PO) that unifies the objective functions in DPO and RLHF. More specifically, for a specific instantiation of $\Psi$, the objective in $\Psi$PO recovers DPO and SLiC-HF. In addition, DPO has been further generalized to include diverse divergence constraints (Wang et al., 2024a). Besides, Hejna et al. (2024) propose contrastive preference learning based on a regret preference model instead of the usual one in RLHF. It is also possible to learn a $Q$ function from human preferences directly, which implies a policy without the need for separate policy- and reward-model training (Myers et al., 2023).
It is worth noting that approaches for directly learning the policy from preferences have been considered in the past as well (Wilson et al., 2012; Fürnkranz et al., 2012; Wirth & Fürnkranz, 2013b; a; Busa-Fekete et al., 2014). In Sections 3.2.1 and 3.2.2 in the survey by Wirth et al. (2017), these methods are explained in more detail.
Another recent trend in fine-tuning models with human feedback is to even manage it without the usage of RL. An alternative is based on supervised reward learning with new types of loss functions (Lee et al., 2023; Yuan et al., 2023) or a specific learning process (Dong et al., 2023; Korbak et al., 2023). There are also RL-free approaches that do not use a reward model to train a policy to execute natural-language instructions using a transformer architecture (Brohan et al., 2023; Yu et al., 2023). On a related note, Liu et al. (2024a) suggest a way how to convert human feedback to natural language sentences for the task of fine-tuning language models.
@@ 7 Theory
The field of RLHF has recently made some progress in terms of theoretical results, which we will discuss in this section. First, we consider the contributions where the goal is to learn a provable (near) optimal policy both in an online and offline fashion or even in a way that falls in between. Then, we discuss and highlight recent contributions related to different theoretical aspects of RLHF, such as its relation to the standard reward-based RL. Tables 6 and 7 provide a concise overview of the results for the online or offline policy learning setting. Here, $\mathcal{N}_{\mathcal{F}}(\epsilon,d)$ denotes the $\epsilon$-covering number of a set $\mathcal{F}$ under some metric $d$ [^4] The $\epsilon$-covering number is the minimum integer $N$ such that there exists a subset $\mathcal{F}^{\prime}\subset\mathcal{F}$ with $|\mathcal{F}^{\prime}|=N$, and for any $f\in\mathcal{F}$, there exists some $f^{\prime}\in\mathcal{F}^{\prime}$ satisfying $d(f,f^{\prime})\leq\epsilon$.. It is worth mentioning that (almost) all works have two standard assumptions, namely that the reward function is bounded and that the ground-truth reward, human feedback model, or transition dynamic is an element of the considered model space, respectively.
@@ 7.1 Policy Learning
In the literature focusing on theoretical results, there is a distinction (similar to the distinction made in standard RL) between an offline and online setting. In the former, learning is based on a given fixed data set, usually previously collected through an interaction with the environment. In contrast, in the online environment, one interacts directly with the environment to learn from real-time feedback and continuously updates one’s strategies based on the feedback received, allowing the agent to learn and adapt as it engages with the environment. Accordingly, an important component of the online variant is the sampling procedure, i.e., how the labels are selected. This is usually accomplished using an acquisition function that is based on uncertainty (see Section 4.1.1).
@@
(Novoseller et al., 2020)
Leveraging Posterior Sampling from dueling bandits Linear link function, tabular MDP Bayes regret minimization w.r.t. optimal policy based on trajectory comparison feedback Asymptotic regret rate:
@@
(Saha et al., 2023)
Leveraging MaxInP from contextual dueling bandits Logistic link function, tabular MDP, linear rewards & $d$-dimensional feature embedding of trajectories Expected regret minimization w.r.t. optimal policy based on trajectory comparison feedback Transition dynamics:
@@ 2. Unknown
(Chen et al., 2022)
Leveraging MaxInP; general function approximation General human feedback model class $\mathcal{F}_{\mathbb{T}}$ and general transition dynamic class $\mathcal{F}_{\mathbb{P}}$ with finite $l_{2}$-norm $\rho$-Eluder dimension $d_{\mathbb{T}}^{(2)}(\rho)$ and $d_{\mathbb{P}}^{(2)}(\rho)$, respectively High probability regret minimization w.r.t. optimal policy based on trajectory comparison feedback $\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{P}}(\tfrac{1}{T})HT\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{P}}}\left(\tfrac{1}{T},d\right)\right)}\right)\newline +\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{T}}(\tfrac{1}{T})T\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{T}}}\left(\tfrac{1}{T},d\right)\right)}\right)$
@@
(Xu et al., 2020)
Dynamic programming, policy search, $(\epsilon,\delta)$-PAC black-box dueling bandit algorithm and simulator Uniform dependence of policy preference probabilities on value function differences, tabular MDP, $(\epsilon,\delta)$-PAC dueling bandit algorithm with $\Psi(K,\varepsilon,\delta)\varepsilon^{-\alpha}$ sample complexity for $K$ arms $(\epsilon,\delta)$-PAC for optimal policy based on trajectory comparison feedback Simulator step bound
@@
(Xu et al., 2020)
Similar to PPS, instead of simulator using an auxiliary synthetic reward function Same as PPS and stochastic triangle inequality of trajectory comparisons preferences $(\epsilon,\delta)$-PAC for optimal policy based on trajectory comparison feedback Step complexity bound
@@
(Kong & Yang, 2022)
Optimistic least-squares value iteration, maximum information gain, value iteration based on pessimistic expected value function estimation Binary rewards for state-action pairs based on human response model $f\in\mathcal{F}_{H}$ with bounded noise $\Delta>0$, compliant and tabular/linear MDP with dimension $d$ $(\epsilon,\delta)$-PAC for optimal policy based on binary state-action reward feedback Tabular MDP:
@@
(Wu & Sun, 2024)
Least-squares value iteration with perturbed state-action-wise reward model General differentiable link function $\Phi$, linear MDP, linear rewards with $d$-dimensional feature embedding of trajectories Expected regret minimization w.r.t. optimal policy and/or low trajectory comparison feedback complexity steered by $\epsilon\in[0,1]$ Expected regret bound:
@@
Algorithm for Policy Alignment in Reinforcement Learning (A-PARL)
(Chakraborty et al., 2024)
Iterative bilevel optimization via gradient descent based on an estimated policy gradient Lipschitz assumptions on the objective function, the reward function, the parametric policy class, and convexity assumptions on the value function Solving the bilevel optimization problem Convergence rate:
$O(1/T)$
@@
The first work dealing with the question of theoretical guarantees for learning an optimal policy from trajectory comparison feedback (see Section 3.2) in an online manner is by Novoseller et al. (2020). It laid the foundation for a paradigm subsequently embraced by many subsequent research endeavors: Adapting learning algorithms from the dueling or preference-based bandit literature (Yue & Joachims, 2009; Sui et al., 2018; Bengs et al., 2021) to the underlying situation with additional states. The preference-based bandit problem can be viewed as a preference-based RL problem with one state, so state transition dynamics must be considered accordingly for a fruitful adaptation. It is worth mentioning that Jain et al. (2015) used a quite similar idea before for feedback in the form of corrections (see Section 3.2) by resorting to the coactive learning setting (Shivaswamy & Joachims, 2012). Assuming the existence of a ground-truth context-trajectory scoring function and that the user’s feedback is informative, the Preference Perceptron algorithm by Shivaswamy & Joachims (2012) is used and analyzed in terms of its cumulative regret.
@@
Novoseller et al. (2020) suggest the Dueling Posterior Sampling (DPS), which is an adaptation of the self-sparring algorithm (Sui et al., 2017). It takes a Bayesian perspective on the problem and defines a Dirichlet prior on the dynamics and a Gaussian prior on the rewards that are subsequently updated, while the trajectories to be compared by the human labeler are chosen based on their (posterior) probability of being optimal [^5] The latter probability is assessed by posterior sampling; a commonly used technique in the bandit literature used by so-called Thompson Sampling strategies, see Lattimore & Szepesvári (2020) for more details.. Assuming a linear link function (see Section 5.1.4) as well as a tabular MDP, it is shown that DPS is (i) consistent, i.e., converges in distribution to the optimal policy, and (ii) achieves an asymptotic expected regret bound (see Table 6).
@@
Xu et al. (2020) combine dynamic programming and policy search with a black-box preference-based bandit algorithm for each state to design routines that return an almost optimal (a.k.a. $\varepsilon$-optimal) policy with high probability [^6] This is a so-called PAC learning setting (Valiant, 1984) in which the goal of finding the/an optimal object is relaxed to finding a “good enough” object, usually specified by some distance measure on the object domain. . The first routine, called Preference-based Policy Search (PPS), requires access to a simulator, while the second routine, called Preference-based Exploration and Policy Search (PEPS), gets rid of this requirement by exploring the state space by means of an auxiliary synthetic reward function. By assuming that the probability of one policy dominating another policy is bounded uniformly over all states from below by a multiplicative of their value function, they show generic upper bounds for both routines on the number of pairwise trajectory comparisons (see Table 6). If these dominance probabilities have even more structural properties, such as fulfilling stochastic transitivity or stochastic triangle inequality (see Haddenhorst et al. (2020); Bengs et al. (2021)), then these upper bounds can be further refined.
@@
A follow-up work by Saha et al. (2023) assumes a feature embedding of trajectories that gives rise to a feature embedding of policies and adapts the MaxInP algorithm (Saha, 2021) for contextual dueling bandits by essentially viewing the policy embeddings as the contexts. More precisely, assuming a logistic link function (see Section 5.1.4), confidence sets for the expected scores of the policies are constructed based on the maximum likelihood estimate (MLE), and the two policies with the highest uncertainty in terms of maximal variance are used to sample a trajectory, respectively, to be compared. In this way, the logistic preference-based reinforcement learning (LPbRL) is derived and also extended to the case of unknown dynamics by taking the uncertainty regarding the dynamics into account when constructing the confidence sets. For both cases, i.e., known or unknown dynamics, upper bounds on the regret of LPbRL are shown (see Table 6).
@@
In contrast to previous work that all considers tabular MDPs, Chen et al. (2022) consider the case of a general unknown human feedback model and unknown dynamics each from function classes with a finite Eluder dimension [^7] Roughly speaking, the Eluder dimension of a function class refers to the number of worst-case errors one must make to identify an unknown function from that class. (Russo & Van Roy, 2013). They propose and analyze the Preference-based Optimistic Planning (PbOP) algorithm, which essentially follows a similar design as LPbRL but uses least-square estimates for the human feedback model and transitions dynamics along with confidence sets based on them. Moreover, they derive lower bounds for the regret of any learning algorithm by reducing the once-per-episode-feedback RL problem (Chatterji et al., 2021) to the PbRL problem. Finally, they extend their analysis to the case of $K$-wise comparisons, where one obtains all $\binom{K}{2}$ pairwise comparisons for $K$ many queried trajectories. In essence, the regret term coming from the human feedback model class improves by a factor of $\sqrt{K}$.